OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. A) y = 2x - 8 B) y = 2x - 4 C) y = 2x + 7 D) y = 2x + 5

OpenStudy (anonymous):

y+2 = 2(x-3) y+2 = 2x-6 y=2x-6-2 y=??

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

I know the problem is -6-2, so let me pose a question, if you go 6 steps backwards and then another 2 backwards, how many steps backwards will you have gone?